A work center has the following input/output table and a standing backlog of 130 hours. What would be backlog be at the end of the third week.

1

2

3

4

5

6

Planned input

300

300

300

300

300

300

Actual Input

305

305

305

305

155

150

Planned output

310

310

310

300

300

300

Actual Output

290

290

290

290

190

150
Answer #1

backlog = previous backlog+planned input - planned output

Planned backlog For period 1 = 130+ 300-310 =120

Actual backlog for period 1 = 130+305-290 =145

Actual backlog for period 2 = 145+305-290 = 160

Actual backlog for period 3 =160+305-290 = 175